OpenStudy (sonyalee77):

HEEEEEEEEEEEEEEEEELP THIS IS SOO HARD! Which of the following conditions would make x + a < x a true statement? A. x equals a B. a is negative C. x is negative D. a is less than x

OpenStudy (tyteen4a03):

B, a is negative. Remember that x + (-a) = x - a.

OpenStudy (sonyalee77):

thanx you helped aot

OpenStudy (anonymous):

@Sonyalee77, do you know how to solve these problems in the future?

OpenStudy (sonyalee77):

no

OpenStudy (anonymous):

You should take the original inequality x + a < x and plug in the conditions. A, B C, and D. Then you can determine whether or not the inquality is true. For example, A is x = a Well, if x = a then, x + a = x + x x + x is NOT less than x, so A cannot be true

OpenStudy (sonyalee77):

thanx for explaining the problem to me

OpenStudy (anonymous):

Can you show me whether or not D is the answer?